OpenStudy (anonymous):

Will GIVE MEDAL and FAN!!(:

OpenStudy (anonymous):

@jim_thompson5910 @ryan123345

OpenStudy (anonymous):

A= 1/2 h (b1 +b2)

jimthompson5910 (jim_thompson5910):

the area of a trapezoid is A = h*(b1 + b2)/2 in this case, h = 7, b1 = 17 and b2 = 13 so A = 7*(17+13)/2 A = ???

jimthompson5910 (jim_thompson5910):

so far so good, just plug in the given info and evaluate

OpenStudy (anonymous):

105? @jim_thompson5910 (:

OpenStudy (anonymous):

The formula for a trapezoid is A=1/2(b1+b2)xh so A= 1/2(17+13)x7 your answer would be 105

OpenStudy (anonymous):

thank U!(:

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Your welcome (=